y^x+x^y+x^x= a^b. Find dy/dx

Grade:10

3 Answers

Roshan Mohanty
10 years ago

dy/dx = [a^b - x^x . ln x . x^x ] / [(x/y) . (y^x)^2 . ln y + (y/x) . (x^y)^2 .ln x]

Anshuman Chandel
10 years ago

put 

u= y^x

v=x^y


w=x^x

use logarithmic differentiation

and dy/dx=du/dx+dv/dx+dw/dx

Gautham Pai M K
10 years ago

Take separately each terms and solve it you get the answer.